1. Consider the reaction

 

The rate of the reaction is when the concentration of A is 0.35 M. Calculate the rate constant if the reaction is (a) first order in A and (b) second order in A.



1
Expert's answer
2021-10-12T07:48:27-0400

First order reaction:


Rate = k(A),




Where (k) is a constant




R = [ ( 1.6)(10-2)](0.35)




R = (5.6)(10-2)ms-1




R = 0.056ms-1




Second order reaction:




Rate = k ( A) 2




R = [(1.6)(10)-2 ](0.35)2




R = (1.96)(10-3)m-1s-1




R = 0.00196m-1s-1